OpenStudy (anonymous):

Solve 2 log x = log 64

myininaya (myininaya):

Using \[\log(x^r)=r \log(x)\] I rewrote the equation \[\log (x^2)=\log (64)\] Set insides equal By the way make sure you don't keep the negative answer if you get one :)

OpenStudy (anonymous):

so it the answer 64?

myininaya (myininaya):

Nope

myininaya (myininaya):

\[x^2=64 \text{ when x=?}\]

OpenStudy (anonymous):

8 !!!

OpenStudy (anonymous):

2logx = log64 logx2 = log64 elogx2=elog64 ##taking exponential both sides## x2=64 x=8

myininaya (myininaya):

yes sofia x=8 is right! :)

OpenStudy (anonymous):

can you help me with this one myininaya, Solve log4 x – 1 = 2

myininaya (myininaya):

\[\log_4(x-1)=2 ?\]

myininaya (myininaya):

or\[\log_4(x)-1=2?\]

myininaya (myininaya):

or neither?

OpenStudy (anonymous):

ayyy i don't know this onee!! :S

myininaya (myininaya):

\[\log(4x-1)=2 \text{ or } \log(4x)-1=2\]

myininaya (myininaya):

I'm just trying to figure out the equation

myininaya (myininaya):

Like is it any of the ones I mentioned?

myininaya (myininaya):

\[\log_4(x)-1=2\] If it is this one which I sorta think is the right interpretation of what you wrote The first step is to add 1 on both sides

myininaya (myininaya):

\[\log_4(x)=3\]

myininaya (myininaya):

Now rewrite as an exponential equation! :) \[x=4^3\]

myininaya (myininaya):

Or I mean an exponential form

myininaya (myininaya):

in*

myininaya (myininaya):

Any questions?